An electronic device designed to operate at any temperature in the range from −6.0°C to 65.4°C contains pure carbon resistors. What is the percent difference in resistance over this range? The temperature coefficient of resistivity for carbon is −0.500 ✕ 10-3/°C.
Answer #1

Let T1 = -6 C

T2 = 65.4 C

alfa = 0.500*10^-3 C

we know, R = Ro + Ro*alfa*(T2 -T1)

R - Ro = Ro*alfa*(T2 - T1)

(R - Ro)/Ro = alfa*(T2 - T1)

the percent difference in resistance over this range = 100*(R - Ro)/Ro

= 100*alfa*(T2 - T1)

= 100*0.5*10^-3*(65.4 - (-6))

= 3.57 %   <<<<<<<-----------Answer